A stock has annual returns of 0.02%, 6.93%, 8.21%, and -2.02% over the past four years. What is the stock’s geometric average return over this period? (Report answer in percentage terms and round to 2 decimal places. Do not round intermediate calculations).
Answer #1

The formula for calculating used to calculate geometric avearge return is:

= ((1+R1)* (1+R2)*(1+R3)…………….(1+Rn))^ 1/n - 1

= ((1+0.002)*(1+0.0693)*(1+0.0821)*(1 - 0.0202))^1/ 4 - 1

= (1.002*1.0693*1.0821*0.9798)^1/4 - 1

= 1.1360 ^0.25 - 1

= 1.0324 - 1

= 0.0324*100

= 3.24%.

In case of any query, kindly comment on the solution.